lexxxxxxxxx:

the supply function for manufacturing a certain item is p(x)=x^2+46x−66 { the demand function is p(x)=56x+30 if x represents the number of items (in hundreds), what is the optimum number of items to be manufactured?

Vocaloid:

optimal # of items occurs when supply = demand so set x^2 + 46x - 66 = 56x + 30 and solve for x then, since each x represents 100 items, you will then multiply the result by 100 let me know if you're having trouble or want to check your solution welcome to QC!

lexxxxxxxxx:

im confused..

lexxxxxxxxx:

i did what you said i go a decimal....

Vocaloid:

hm. x^2 + 46x - 66 = 56x + 30 start by combining like terms. first, subtract 56x from both sides.

lexxxxxxxxx:

i did

Vocaloid:

good so x^2 - 10x - 66 = 30 now try subtracting 30 from each side

lexxxxxxxxx:

i thought you add 66 to both sides :(

lexxxxxxxxx:

but okay

Vocaloid:

you could add 66 if you wanted to, as long as all the terms end up on one side adding 66 would give us x^2 - 10x = 96 then you can subtract 96 from both sides to get x^2 - 10x - 96 now, you will need to find two numbers that multiply to -96 and add up to -10 lmk if you need some help with this

lexxxxxxxxx:

-16+6

Vocaloid:

good so it factors to (x-16)(x+6) = 0 what is x equal to? keep in mind you only need to consider the positive x solution here

lexxxxxxxxx:

i dont even know... tbh

Vocaloid:

if you have A*B = 0 you know that either A = 0 or B = 0 so (x-16) = 0 (x+6) = 0 solve for x, you will get one positive and one negative value

lexxxxxxxxx:

-6+6=0 ; 16-16=0

Vocaloid:

good so x - 16 = 0 means that x = 16 x + 6 = 0 means that x = -6 since x = the number of items, this cannot be negative so we only consider x = 16 since each x represents 100 items, multiply this by 100 to get 1600 = your solution

lexxxxxxxxx:

thanks so much !